OpenStudy (sanchez9457):

Does the series (n/(ln n)^2) converge or diverege? when i do the nth term test, it doesn't tell me anything but when i do the limit comparison test, it diverges, why is this? Should the nth term test show divergence if the limit comparison test shows it?

OpenStudy (anonymous):

Not all tests work for all series, which is one of the reasons there are so many tests. The only way you'd really have a problem is if one test said it diverged and another said it converged.

OpenStudy (anonymous):

if \( \displaystyle \lim_{n \rightarrow +\infty} u_n \neq 0\) then \(\displaystyle \sum_{n=1}^{+\infty} u_n\) is divergent